Circular No. 6777

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA IN NGC 205 Q. Y. Qiao, Y. L. Qiu, J. Y. Hu, and W. D. Li, Beijing Astronomical

Observatory (BAO); A. Esamdin, Urumqi Astronomical Station; and Y. Zhang, Beijing Normal University, report their discovery of a nova in NGC 205 on Nov. 21.54 UT, in the course of the BAO supernova survey, using the 0.6-m reflector at Xinglong Station. The nova is at α = 0^{h}40^{m}23^{s}.26, δ = +41^{o}43′27′′.0 (equinox 2000.0), which is 13′′.2 east and 139′′.9 north of the nucleus of NGC 205. Unfiltered CCD magnitudes: Nov. 21.54, 18.8; 22.50, 18.4; 23.44, 18.5. The nova was not present on a CCD image taken on Nov. 4. A low-resolution spectrogram of the nova was obtained with the 2.16-m telescope at Xinglong on Nov. 23.4 by J. Y. Wei, L. Cao (BAO), and Q. S. Gu (Nanjing University); the spectrum shows typical features of a nova after maximum. The Balmer emission lines are very strong (equivalent width of Hβ is −142.4 km/s). There are also strong Fe II emission lines (492.4, 451.5, 458.4, 516.9, 527.6, 532.6, and 590.1 nm), He I emission lines (492.2 and 501.6 nm), and O I emission lines (615.8 and 777.6 nm). The emission lines show blueshift with a velocity of 224 km/s, which is similar to that of NGC 205. The V magnitude derived from the flux at 550.0 nm continuum is 18.1, yielding an absolute V magnitude of −6.6 (24.7 adopted as the distance modulus of NGC 205).

SMC X-3 F. E. Marshall, Goddard Space Flight Center (GSFC); and J. C.

Lochner and T. Takeshima, GSFC and Universities Space Research Association, report the detection with the RXTE PCA of an outburst from the x-ray transient SMC X-3 and the discovery of a period of 92 ± 1.5 s with a complex pulse profile: “First detected on Nov. 15.44 UT with a 2–10-keV flux of 1.5 mCrab, the source flux had increased to ∼ 4 mCrab by Nov. 25.44. Further RXTE observations are planned, and observations of the suggested optical counterpart are encouraged.”

SUPERNOVA 1997dp IN NGC 969 Unfiltered CCD magnitude estimate of SN 1997dp by S. Moretti and

S. Tomaselli, Forlí, Italy: Nov. 3.89 UT, 17.9 (GSC reference stars).

1997 November 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green